# Variational Diffusion Posterior Sampling with Midpoint Guidance

### v1

@@ 3 The MGPS algorithm
4:     $\hat{\mu}_{{\ell_{k}}}\leftarrow\frac{\sqrt{\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}}\smash{m^{\theta}_{0|{\ell_{k+1}}}}(\hat{X}_{{\ell_{k+1}}})+\frac{\sqrt{\alpha_{k+1}/\alpha_{{\ell_{k}}}}(1-\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}}X_{k+1}$
5:     $\hat{\rho}_{\ell_{k}}\leftarrow\frac{1}{2}\log\frac{(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})(1-\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}}$
@@
7:        $Z\sim\mathcal{N}(\mathrm{0}_{d},\mathrm{I}_{d})$
8:        $(\hat{\mu}_{\ell_{k}},\hat{\rho}_{\ell_{k}})\leftarrow\mathsf{OptimizerStep}(\nabla_{\varphi}\widetilde{\mathcal{L}}_{k}(\cdot,Z;X_{k+1});\,\hat{\mu}_{\ell_{k}},\hat{\rho}_{\ell_{k}})$
@@
10:     $Z_{\ell_{k}},Z_{k}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{N}(\mathrm{0}_{d},\mathrm{I}_{d})$
11:     $\hat{X}_{\ell_{k}}\leftarrow\hat{\mu}_{\ell_{k}}+\mathrm{diag}(\mathrm{e}^{\hat{\rho}_{\ell_{k}}})Z_{\ell_{k}}$
12:     $X_{k}\leftarrow\frac{\sqrt{\alpha_{k}/\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{k})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}\hat{X}_{\ell_{k}}+\frac{\sqrt{\alpha_{k+1}/\alpha_{k}}(1-\alpha_{k}/\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}X_{k+1}+\frac{\sqrt{(1-\alpha_{k+1}/\alpha_{k})(1-\alpha_{k}/\alpha_{{\ell_{k}}})}}{\sqrt{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}}Z_{k}$
@@
14:  $X_{0}\leftarrow\smash{m^{\theta}_{0|1}}(X_{1})$
We now provide a summary of the MGPS algorithm, whose pseudocode is given in Algorithm 1. Given $({\ell_{k}})_{k=1}^{n-1}$, MGPS proceeds by simulating a Markov chain $(X_{k})_{k=n}^{0}$ starting from $X_{n}\sim\mathcal{N}(\mathrm{0}_{d},\mathrm{I}_{d})$. Recursively, given the state $X_{k+1}$, the state $X_{k}$ is obtained by
@@
    minimizing $\mathcal{L}_{k}(\cdot;X_{k+1})$ by performing $M$ stochastic gradient steps using the gradient estimator $\nabla_{\varphi}\widetilde{\mathcal{L}}_{k}(\cdot;X_{k+1})$, yielding a parameter $\varphi^{*}_{k}(X_{k+1})$,
@@
    sampling $\hat{X}_{\ell_{k}}\sim\lambda^{\varphi^{*}_{k}}_{{\ell_{k}}|k+1}$, where we drop the dependence on $X_{k+1}$ in $\varphi^{*}_{k}$, and then $X_{k}\sim q_{k|{\ell_{k}},k+1}(\cdot|\hat{X}_{\ell_{k}},X_{k+1})$.
@@
While conditionally on $\varphi^{*}_{k}$ we have that $\smash{\hat{X}_{\ell_{k}}\sim\lambda^{\varphi^{*}_{k}}_{{\ell_{k}}|k+1}}$, the actual law of $\hat{X}_{\ell_{k}}$ given $X_{k+1}$ is not a Gaussian distribution as we need to marginalize over that of $\varphi^{*}_{k}$ due to the randomness in the stochastic gradients.
A well-chosen initialization for the Gaussian variational approximation parameters is crucial for achieving accurate fitting with few optimization steps, ensuring the overall runtime of MGPS remains competitive with existing methods. Given $\hat{X}_{\ell_{k+1}}$ sampled from $\smash{\lambda^{\varphi^{*}_{k+1}}_{{\ell_{k+1}}|k+2}}$ at the previous step, we choose the initial parameter $\varphi_{k}$ so that
$\lambda^{\varphi_{k}}_{{\ell_{k}}|k+1}(x_{k})=\smash{q_{{\ell_{k}}|0,k+1}(x_{k}|\smash{m^{\theta}_{0|{\ell_{k+1}}}}(\hat{X}_{\ell_{k+1}}),X_{k+1})}\,.$
This initialization is motivated by noting that $q_{{\ell_{k}}|0,k+1}(\cdot|m^{\pi}_{0|k+1}(X_{k+1}),X_{k+1})$, the DDPM approximation of $\smash{\pi}_{{\ell_{k}}|k+1}(\cdot|X_{k+1})$, where $m^{\pi}_{0|k+1}$ is a supposed denoiser for the posterior, is a reasonable candidate for the initialization. As it is intractable, we replace it with $\smash{m^{\theta}_{0|{\ell_{k+1}}}}(\hat{X}_{\ell_{k+1}})$, our best current guess.
Finally, in Section C.2 we devise a warm-start approach that improves the initialization during the early steps of the algorithm. This has proved advantageous in the most challenging problems.
@@
When $g$ is a likelihood associated to a linear inverse problem with Gaussian noise, Finzi et al. (2023); Stevens et al. (2023); Boys et al. (2023) leverage the fact that the Gaussian projection of $\smash{q_{0|k}(\cdot|x_{k})}$ minimizing the forward KL divergence can be approximated using the pre-trained denoisers; see Meng et al. (2021, Theorem 1). In the considered Gaussian setting, the potential $g$ can be exactly integrated against the estimated Gaussian projection. However, this involves a matrix inversion that may be prohibitively expensive. Boys et al. (2023) circumvent the latter by using a diagonal approximation which involves the computation of ${d}$ vector-Jacobian products. For general likelihoods $g$, Song et al. (2023b) use the Gaussian approximations of Ho et al. (2022); Song et al. (2023a) to estimate $g_{k}$ for non-linear potentials $g$ using a vanilla Monte Carlo estimate. Zhang et al. (2023) also consider the surrogate transitions (3.5) with ${\ell_{k}}=k$, and, given an approximate sample $X_{k+1}$ from $\hat{\pi}^{\ell}_{k+1}$, estimate the next sample $X_{k}$ maximizing $x_{k}\mapsto\smash{\hat{\pi}^{\theta}_{{k}|k+1}(x_{k}|X_{k+1})}$ using gradient ascent. They also consider improved estimates of the potential $g_{k}$ running a few diffusion steps. However, it is shown (Zhang et al., 2023, last subtable in Table 2) that this brings minor improvements at the expense of a sharp increase in computational cost, which is mainly due to backpropagation over the denoisers.
Finally, a second line of work considers the distribution path $(\hat{\pi}_{k})_{k=n}^{0}$, where $\hat{\pi}_{k}(x_{k})\propto g(\smash{m_{0|k}}(x_{k}))q_{k}(x_{k})$ for $k\in\llbracket 0,n-1\rrbracket$ and $\hat{\pi}_{n}=\mathcal{N}(\mathrm{0}_{d},\mathrm{I}_{d})$. This path bridges the Gaussian distribution and the posterior of interest. Furthermore, if one is able to accurately sample from $\hat{\pi}_{k+1}$, then these samples can be used to initialize a sampler targeting the next distribution $\hat{\pi}_{k}$. As these distributions are expected to be close, the sampling from $\hat{\pi}_{k}$ can also be expected to be accurate. Repeating this process yields approximate samples from $\pi$ regardless of the approximation error in the potentials. This approach is pursued by Rozet & Louppe (2023), who combines the update (2.5) with a Langevin dynamics targeting $\hat{\pi}_{k}$. Wu et al. (2023) use instead sequential Monte Carlo Gordon et al. (1993) to recursively build empirical approximations of each $\hat{\pi}_{k}$ by evolving a sample of $N$ particles.
@@ 4 Experiments
We now evaluate our algorithm on three different problems and compare it with several competitors. We begin in benchmarking our method on toy Gaussian-mixture targets and image experiments with both pixel space and latent diffusion models. We review the latter in Section A.2. Finally, we perform inpainting experiments on ECG data. For experiments based on pixel-space diffusion models, MGPS is benchmarked against state-of-the art methods in the literature: DPS Chung et al. (2023), PGDM Song et al. (2023a), DDNM Wang et al. (2023), DiffPir Zhu et al. (2023), and RedDiff Mardani et al. (2024). Regarding experiments using latent diffusion models, we compare against PSLD Rout et al. (2024) and ReSample Song et al. (2024). Full details on the hyperparameters can be found in Section D.1.
@@
We first evaluate the accuracy of our method and the impact of the hyperparameters on a toy linear inverse problem with Gaussian mixture (GM) as $q$ and a Gaussian likelihood $g(x)\vcentcolon=\mathcal{N}(y;\mathrm{A}x,\sigma_{y}^{2}\mathrm{I}_{d})$, where $\mathrm{A}\in\mathbb{R}^{{d_{y}}\times{d}}$ and $\sigma_{y}=0.05$. We repeat the experiment of Cardoso et al. (2023, App B.3), where the prior is a GM with 25 well-separated strata. For this specific example, both the posterior and the denoiser $\smash{m_{0|k}}$ are available in a closed form. In particular, the posterior can be shown to be a GM. The full details are provided in Section D.3. We consider the settings $(20,1)$ and $(200,1)$ for the dimensions $({d},{d_{y}})$. For each method, we reconstruct $1000$ samples and then compute the sliced Wasserstein (SW) distance to exact samples from the posterior distribution. We repeat this procedure $100$ times with randomly drawn matrices $\mathrm{A}$, and consider the resulting averaged SW distance. In every replication, the observation is generated as $Y=\mathrm{A}X+\sigma_{y}Z$ where $Z\sim\mathcal{N}(\mathrm{0}_{d_{y}},\mathrm{I}_{d_{y}})$ and $X\sim q$. The results are reported in Section 4. It can be observed that MGPS with $\eta=0.75$ outperforms all baselines in both settings. Although we have tuned the parameters of DPS  it still exhibits considerable instability and often diverges. To account for these instabilities we set the SW to $10$ when it diverges.
@@
Table 1: 95 % confidence interval for the SW on the GM experiment. ${d}=20,{d_{y}}=1$ ${d}=200,{d_{y}}=1$ MGPS ${\bf{2.11}}\pm 0.30$ ${\bf{3.66}}\pm 0.53$ DPS $8.93\pm 0.49$ $9.15\pm 0.44$ PGDM $\underline{2.44}\pm 0.36$ $\underline{5.23}\pm 0.38$ DDNM $4.24\pm 0.37$ $7.10\pm 0.50$ DiffPir $4.14\pm 0.42$ $8.43\pm 0.92$ RedDiff $6.70\pm 0.45$ $8.35\pm 0.39$
@@
We evaluate our algorithm on a wide range of linear and nonlinear inverse problems with noise level $\sigma_{y}=0.05$. As linear problems we consider: image inpainting with a box mask of shape $150\times 150$ covering the center of the image and half mask covering its right-hand side; image Super Resolution (SR) with factors ×$\times 4$ and ×$\times 16$; image deblurring with Gaussian blur; motion blur with $61\times 61$ kernel size. We use the same experimental setup as Chung et al. (2023, Section 4) for the last two tasks. Regarding the nonlinear inverse problems on which we benchmark our algorithm, we consider: JPEG dequantization with quality factor $2\%$; phase retrieval with oversampling factor $2$; non-uniform deblurring emulated via the forward model of Tran et al. (2021); high dynamic range (HDR) following Mardani et al. (2024, Section 5.2). Since the JPEG operator is not differentiable, we instead use the differentiable JPEG framework from Shin & Song (2017). We note that only DPS and RedDiff apply to nonlinear problems.
Datasets and models. We test our algorithm on the $256\times 256$ versions of the FFHQ Karras et al. (2019) and ImageNet Deng et al. (2009) datasets. We chose randomly $50$ images from each to be used in the evaluation of the algorithms. We leverage pre-trained DDMs that are publicly available. For FFHQ, we use the pixel-space DDM of Choi et al. (2021) and the LDM of Rombach et al. (2022) with VQ4 autoencoder. We use the model by Dhariwal & Nichol (2021) for ImageNet.
@@
Evaluation. We use LPIPS Zhang et al. (2018) as the metric to evaluate the quality of the reconstructions. Given that the posterior distribution is expected to be highly multimodal for most of the tasks, pixel-wise comparisons with the ground truth, such as PSNR and SSIM, provide limited insight into reconstruction quality. For all the considered algorithms, we use the hyperparameters proposed in their official implementations if available, otherwise we manually tune them to achieve the best reconstructions. We defer implementation details to Section D.1. For each inverse problem, we compute the LPIPS between the ground-truth image and a reconstruction and then average over the dataset. For phase retrieval, however, due to the task’s inherent instability, we follow the approach of Chung et al. (2023) by selecting the best reconstruction out of four; further discussion on this can be found in Chung et al. (2023, Appendix C.6).
@@
Results.  We report the results in Table 2 for FFHQ and ImageNet with DDM prior, and Section 4 for FFHQ with LDM prior. These reveal that MGPS consistently outperforms the competing algorithms across both linear and nonlinear problems. Notably, on nonlinear tasks, MGPS achieves up to a twofold reduction in LPIPS compared to the other methods with DDM prior. It effectively manages the additional nonlinearities introduced by using LDMs and surpasses the state-of-the-art, as demonstrated in Section 4. Reconstructions obtained with MGPS are displayed in Figure 2 and Figure 4. Further examples and comparisons with other competitors are provided in Section D.6. It is seen that MGPS provides high quality reconstructions even for the most challenging examples. For the DDM prior, the results are obtained using the warm-start approach, see Algorithm 3, and setting ${\ell_{k}}=\lfloor k/2\rfloor$ on all the tasks based on the FFHQ  dataset. As for the ImageNet dataset, we use the same configuration on all the tasks, except for Gaussian deblur and motion deblur. For these tasks we found that using ${\ell_{k}}=\lfloor k/2\rfloor\mathbbm{1}_{k\geq\lfloor n/2\rfloor}+k\mathbbm{1}_{k<\lfloor n/2\rfloor}$ improves performance. We use a similar strategy for the LDM prior. In Section D.2.1, we measure the runtime and GPU memory requirements for each algorithm. The memory requirement of our algorithm is the same as that of DPS and PGDM, but the runtime with $n=300$ is larger. With fewer diffusion steps, the runtime is halved compared to DPS and PGDM and comparable to that of DiffPir, DDNM and RedDiff. Still, MGPS remains competitive and consistently outperforms the baselines, particularly on nonlinear tasks. See Table 7 and 8 for detailed results using $n\in\{50,100\}$ diffusion steps.
@@
Table 2: Mean LPIPS for various linear and nonlinear imaging tasks on the FFHQ and ImageNet  $256\times 256$ datasets with $\sigma_{y}=0.05$.

### v2

@@ 3 The MGPS algorithm
4:     $\hat{\boldsymbol{\mu}}_{{\ell_{k}}}\leftarrow\frac{\sqrt{\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}}\smash{\boldsymbol{m}^{\theta}_{0|{\ell_{k+1}}}}(\hat{\mathbf{X}}_{{\ell_{k+1}}})+\frac{\sqrt{\alpha_{k+1}/\alpha_{{\ell_{k}}}}(1-\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}}\mathbf{X}_{k+1}$
5:     $\hat{\boldsymbol{\rho}}_{\ell_{k}}\leftarrow\frac{1}{2}\log\frac{(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})(1-\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}}$
@@
7:        $\mathbf{Z}\sim\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d})$
8:        $(\hat{\boldsymbol{\mu}}_{\ell_{k}},\hat{\boldsymbol{\rho}}_{\ell_{k}})\leftarrow\mathsf{OptimizerStep}(\nabla_{\boldsymbol{\varphi}}\widetilde{\mathcal{L}}_{k}(\cdot,\mathbf{Z};\mathbf{X}_{k+1});\,\hat{\boldsymbol{\mu}}_{\ell_{k}},\hat{\boldsymbol{\rho}}_{\ell_{k}})$
@@
10:     $\mathbf{Z}_{\ell_{k}},\mathbf{Z}_{k}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d})$
11:     $\hat{\mathbf{X}}_{\ell_{k}}\leftarrow\hat{\boldsymbol{\mu}}_{\ell_{k}}+\mbox{diag}(\mathrm{e}^{\hat{\boldsymbol{\rho}}_{\ell_{k}}})\mathbf{Z}_{\ell_{k}}$
12:     $\mathbf{X}_{k}\sim q_{k|{\ell_{k}},k+1}(\cdot|\hat{\mathbf{X}}_{\ell_{k}},\mathbf{X}_{k+1})$ (See (A.1))
@@
14:  $\mathbf{X}_{0}\leftarrow\smash{\boldsymbol{m}^{\theta}_{0|1}}(\mathbf{X}_{1})$
We now provide a summary of the MGPS algorithm, whose pseudocode is given in Algorithm 1. Given $({\ell_{k}})_{k=1}^{n-1}$, MGPS proceeds by simulating a Markov chain $(\mathbf{X}_{k})_{k=n}^{0}$ starting from $\mathbf{X}_{n}\sim\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d})$. Recursively, given the state $\mathbf{X}_{k+1}$, the state $\mathbf{X}_{k}$ is obtained by
@@
    minimizing $\mathcal{L}_{k}(\cdot;\mathbf{X}_{k+1})$ by performing $M$ stochastic gradient steps using the gradient estimator $\nabla_{\boldsymbol{\varphi}}\widetilde{\mathcal{L}}_{k}(\cdot;\mathbf{X}_{k+1})$, yielding a parameter ${\boldsymbol{\varphi}}^{*}_{k}(\mathbf{X}_{k+1})$,
@@
    sampling $\hat{\mathbf{X}}_{\ell_{k}}\sim\lambda^{{\boldsymbol{\varphi}}^{*}_{k}}_{{\ell_{k}}|k+1}$, where we drop the dependence on $\mathbf{X}_{k+1}$ in ${\boldsymbol{\varphi}}^{*}_{k}$, and then $\mathbf{X}_{k}\sim q_{k|{\ell_{k}},k+1}(\cdot|\hat{\mathbf{X}}_{\ell_{k}},\mathbf{X}_{k+1})$.
@@
While conditionally on ${\boldsymbol{\varphi}}^{*}_{k}$ we have that $\smash{\hat{\mathbf{X}}_{\ell_{k}}\sim\lambda^{{\boldsymbol{\varphi}}^{*}_{k}}_{{\ell_{k}}|k+1}}$, the actual law of $\hat{\mathbf{X}}_{\ell_{k}}$ given $\mathbf{X}_{k+1}$ is not a Gaussian distribution as we need to marginalize over that of ${\boldsymbol{\varphi}}^{*}_{k}$ due to the randomness in the stochastic gradients.
A well-chosen initialization for the Gaussian variational approximation parameters is crucial for achieving accurate fitting with few optimization steps, ensuring the overall runtime of MGPS remains competitive with existing methods. Given $\hat{\mathbf{X}}_{\ell_{k+1}}$ sampled from $\smash{\lambda^{{\boldsymbol{\varphi}}^{*}_{k+1}}_{{\ell_{k+1}}|k+2}}$ at the previous step, we choose the initial parameter ${\boldsymbol{\varphi}}_{k}$ so that
$\lambda^{{\boldsymbol{\varphi}}_{k}}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{k})=\smash{q_{{\ell_{k}}|0,k+1}(\boldsymbol{x}_{k}|\smash{\boldsymbol{m}^{\theta}_{0|{\ell_{k+1}}}}(\hat{\mathbf{X}}_{\ell_{k+1}}),\mathbf{X}_{k+1})}\,.$ (3.9)
This initialization is motivated by noting that $q_{{\ell_{k}}|0,k+1}(\cdot|\boldsymbol{m}^{\pi}_{0|k+1}(\mathbf{X}_{k+1}),\mathbf{X}_{k+1})$, the DDPM approximation of $\smash{\pi}_{{\ell_{k}}|k+1}(\cdot|\mathbf{X}_{k+1})$, where $\boldsymbol{m}^{\pi}_{0|k+1}$ is a supposed denoiser for the posterior, is a reasonable candidate for the initialization. As it is intractable, we replace it with $\smash{\boldsymbol{m}^{\theta}_{0|{\ell_{k+1}}}}(\hat{\mathbf{X}}_{\ell_{k+1}})$, our best current guess. Finally, in Section C.2 we devise a warm-start approach that improves the initialization during the early steps of the algorithm. This has proved advantageous in the most challenging problems.
@@
When $p({\boldsymbol{y}}|\cdot)$ is a likelihood associated to a linear inverse problem with Gaussian noise, Finzi et al. (2023); Stevens et al. (2023); Boys et al. (2023) leverage the fact that the Gaussian projection of $\smash{q_{0|k}(\cdot|\boldsymbol{x}_{k})}$ minimizing the forward KL divergence can be approximated using the pre-trained denoisers; see Meng et al. (2021, Theorem 1). In the considered Gaussian setting, the likelihood $p({\boldsymbol{y}}|\cdot)$ can be exactly integrated against the estimated Gaussian projection. However, this involves a matrix inversion that may be prohibitively expensive. Boys et al. (2023) circumvent the latter by using a diagonal approximation which involves the computation of ${d}$ vector-Jacobian products. For general likelihoods $p({\boldsymbol{y}}|\cdot)$, Song et al. (2023b) use the Gaussian approximations of Ho et al. (2022); Song et al. (2023a) to estimate $p_{k}({\boldsymbol{y}}|\cdot)$ for non-linear likelihoods $p({\boldsymbol{y}}|\cdot)$ using a vanilla Monte Carlo estimate. Zhang et al. (2023) also consider the surrogate transitions (3.5) with ${\ell_{k}}=k$, and, given an approximate sample $\mathbf{X}_{k+1}$ from $\hat{\pi}^{\ell}_{k+1}$, estimate the next sample $\mathbf{X}_{k}$ maximizing $\boldsymbol{x}_{k}\mapsto\smash{\hat{\pi}^{\theta}_{{k}|k+1}(\boldsymbol{x}_{k}|\mathbf{X}_{k+1})}$ using gradient ascent. They also consider improved estimates of the likelihood $p_{k}({\boldsymbol{y}}|\cdot)$ running a few diffusion steps. However, it is shown (Zhang et al., 2023, last subtable in Table 2) that this brings minor improvements at the expense of a sharp increase in computational cost, which is mainly due to backpropagation over the denoisers. See also Section C.3 for a discussion on how our method relates to the DPS Chung et al. (2023) in the case where ${\ell_{k}}=k$.
Finally, a second line of work considers the distribution path $(\hat{\pi}_{k})_{k=n}^{0}$, where $\hat{\pi}_{k}(\boldsymbol{x}_{k})\propto p({\boldsymbol{y}}|\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k}))q_{k}(\boldsymbol{x}_{k})$ for $k\in\llbracket 0,n-1\rrbracket$ and $\hat{\pi}_{n}=\mathcal{N}(\mathbf{0}_{d},\mathbf{I}_{d})$. This path bridges the Gaussian distribution and the posterior of interest. Furthermore, if one is able to accurately sample from $\hat{\pi}_{k+1}$, then these samples can be used to initialize a sampler targeting the next distribution $\hat{\pi}_{k}$. As these distributions are expected to be close, the sampling from $\hat{\pi}_{k}$ can also be expected to be accurate. Repeating this process yields approximate samples from $\pi$ regardless of the approximation error in the likelihoods. This approach is pursued by Rozet & Louppe (2023), who combines the update (2.5) with a Langevin dynamics targeting $\hat{\pi}_{k}$. Wu et al. (2023) use instead sequential Monte Carlo Gordon et al. (1993) to recursively build empirical approximations of each $\hat{\pi}_{k}$ by evolving a sample of $N$ particles.
@@ 4 Experiments
We now evaluate our algorithm on three different problems and compare it with several competitors. We begin in benchmarking our method on toy Gaussian-mixture targets and image experiments with both pixel space and latent diffusion models. We review the latter in Section A.2. Finally, we perform inpainting experiments on ECG data. For experiments based on pixel-space diffusion models, MGPS is benchmarked against state-of-the art methods in the literature: DPS Chung et al. (2023), PGDM Song et al. (2023a), DDNM Wang et al. (2023), DiffPir Zhu et al. (2023), and RedDiff Mardani et al. (2024). Regarding experiments using latent diffusion models, we compare against PSLD Rout et al. (2024) and ReSample Song et al. (2024). Full details on the hyperparameters can be found in Section D.1. Our code for reproducing all the experiments is publicly available. [^1] Code available at https://github.com/yazidjanati/mgps
@@
We first evaluate the accuracy of our method and the impact of the hyperparameters on a toy linear inverse problem with Gaussian mixture (GM) as $q$ and a Gaussian likelihood $p({\boldsymbol{y}}|\boldsymbol{x})\vcentcolon=\mathcal{N}({\boldsymbol{y}};\mathbf{A}\boldsymbol{x},\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d})$, where $\mathbf{A}\in\mathbb{R}^{{d_{\boldsymbol{y}}}\times{d}}$ and $\sigma_{\boldsymbol{y}}=0.05$. We repeat the experiment of Cardoso et al. (2023, App B.3), where the prior is a GM with 25 well-separated strata. For this specific example, both the posterior and the denoiser $\smash{\boldsymbol{m}_{0|k}}$ are available in a closed form. In particular, the posterior can be shown to be a GM. The full details are provided in Section D.3. We consider the settings $(20,1)$ and $(200,1)$ for the dimensions $({d},{d_{\boldsymbol{y}}})$. For each method, we reconstruct $1000$ samples and then compute the sliced Wasserstein (SW) distance to exact samples from the posterior distribution. We repeat this procedure $100$ times with randomly drawn matrices $\mathbf{A}$, and consider the resulting averaged SW distance. In every replication, the observation is generated as $\mathbf{Y}=\mathbf{A}\mathbf{X}+\sigma_{\boldsymbol{y}}\mathbf{Z}$ where $\mathbf{Z}\sim\mathcal{N}(\mathbf{0}_{d_{\boldsymbol{y}}},\mathbf{I}_{d_{\boldsymbol{y}}})$ and $\mathbf{X}\sim q$. The results are reported in Section 4. It can be observed that MGPS with $\eta=0.75$ outperforms all baselines in both settings. Although we have tuned the parameters of DPS  it still exhibits considerable instability and often diverges. To account for these instabilities we set the SW to $10$ when it diverges.
@@
Table 1: 95 % confidence interval for the SW on the GM experiment. ${d}=20,{d_{\boldsymbol{y}}}=1$ ${d}=200,{d_{\boldsymbol{y}}}=1$ MGPS ${\bf{2.11}}\pm 0.30$ ${\bf{3.66}}\pm 0.53$ DPS $8.93\pm 0.49$ $9.15\pm 0.44$ PGDM $\underline{2.44}\pm 0.36$ $\underline{5.23}\pm 0.38$ DDNM $4.24\pm 0.37$ $7.10\pm 0.50$ DiffPir $4.14\pm 0.42$ $8.43\pm 0.92$ RedDiff $6.70\pm 0.45$ $8.35\pm 0.39$
@@
We evaluate our algorithm on a wide range of linear and nonlinear inverse problems with noise level $\sigma_{\boldsymbol{y}}=0.05$. As linear problems we consider: image inpainting with a box mask of shape $150\times 150$ covering the center of the image and half mask covering its right-hand side; image Super Resolution (SR) with factors ×$\times 4$ and ×$\times 16$; image deblurring with Gaussian blur; motion blur with $61\times 61$ kernel size. We use the same experimental setup as Chung et al. (2023, Section 4) for the last two tasks. Regarding the nonlinear inverse problems on which we benchmark our algorithm, we consider: JPEG dequantization with quality factor $2\%$; phase retrieval with oversampling factor $2$; non-uniform deblurring emulated via the forward model of Tran et al. (2021); high dynamic range (HDR) following Mardani et al. (2024, Section 5.2). Since the JPEG operator is not differentiable, we instead use the differentiable JPEG framework from Shin & Song (2017). We note that only DPS and RedDiff apply to nonlinear problems.
@@
Datasets and evaluation. We test our algorithm on the $256\times 256$ versions of the FFHQ Karras et al. (2019) and ImageNet Deng et al. (2009) datasets using publicly available pre-trained DDMs. For FFHQ, we use the pixel-space DDM of Choi et al. (2021) and the LDM of Rombach et al. (2022) with the VQ4 autoencoder. For ImageNet, we use the model by Dhariwal & Nichol (2021). Due to computational constraints, we first evaluate MGPS and competitors on all tasks using 50 random images per dataset. We then focus on the 5 most challenging tasks, evaluating MGPS and the top 2 competitors on 1k images in Section D.4. For all tasks, we report the LPIPS Zhang et al. (2018) between the reference image and the reconstruction, averaged over 50 images. For the five tasks tested on 1k images, we also report the FID in Section D.4. Although pixel-wise metrics like PSNR and SSIM are less informative for multimodal posterior distributions, they are available in Sections D.4, 8, 9 and D.4. For all the considered competitors, we use the hyperparameters proposed in their official implementations if available, otherwise we manually tune them to achieve the best reconstructions (implementation details in Section D.1). For phase retrieval, due to the task’s inherent instability, we follow the approach of Chung et al. (2023) by selecting the best reconstruction out of four; further discussion on this can be found in Chung et al. (2023, Appendix C.6).
@@
Results.  We report the results in Table 2 for FFHQ and ImageNet with DDM prior, and Section 4 for FFHQ with LDM prior. These reveal that MGPS consistently outperforms the competing algorithms across both linear and nonlinear problems. Notably, on nonlinear tasks, MGPS achieves up to a twofold reduction in LPIPS compared to the other methods with DDM prior. It effectively manages the additional nonlinearities introduced by using LDMs and surpasses the state-of-the-art, as demonstrated in Section 4. Reconstructions obtained with MGPS are displayed in Figure 3 and Figure 5. Further examples and comparisons with other competitors are provided in Section D.6. It is seen that MGPS provides high quality reconstructions even for the most challenging examples. For the DDM prior, the results are obtained using the warm-start approach, see Algorithm 3, and setting ${\ell_{k}}=\lfloor k/2\rfloor$ on all the tasks based on the FFHQ  dataset. As for the ImageNet dataset, we use the same configuration on all the tasks, except for Gaussian deblur and motion deblur. For these tasks we found that using ${\ell_{k}}=\lfloor k/2\rfloor\mathbbm{1}_{k\geq\lfloor n/2\rfloor}+k\mathbbm{1}_{k<\lfloor n/2\rfloor}$ improves performance. We use a similar strategy for the LDM prior. In Section D.2.1, we measure the runtime and GPU memory requirements for each algorithm. The memory requirement of our algorithm is the same as that of DPS and PGDM, but the runtime of MGPS with $n=300$ is slightly larger. With fewer diffusion steps, the runtime is halved compared to DPS and PGDM and comparable to that of DiffPir, DDNM and RedDiff. Still, MGPS remains competitive and consistently outperforms the baselines, particularly on nonlinear tasks. See Table 8 and 9 for detailed results using $n\in\{50,100\}$ diffusion steps.
@@
Table 2: Mean LPIPS for various linear and nonlinear imaging tasks on the FFHQ and ImageNet  $256\times 256$ datasets with $\sigma_{\boldsymbol{y}}=0.05$.